OpenStudy (anonymous):

Devante paid a security deposit of $525 when he signed the lease for his apartment. The money was put into an account earning 2% simple annual interest. After renting for 8 years, Devante got his deposit back. What was the payout balance?

OpenStudy (johnweldon1993):

So the formula for simple interest...is \[I = prt\] I = interest earned (what we want) p = original balance r = rate (decimal) t = time (in years) so after plugging in all the numbers we have...we get \[\large I = 525 \times .02 \times 8\] This will solve for the amount of interest you have earned....you then add this to the original 525 to get the answer you want... Hope that helps @rsebren123